how much water must be added to 12 grams of 90% iodine solution to produce a 25% iodine solution?
Answer by checkley71(8403)

THE ORIGINAL SOLUTION CONTAINED 10% WATER TO WHICH X GRAMS OF 100% WATER IS ADDED TO OBTAIN A MIX OR 75% WATER THUS
12*.1+X=.75(12+X) OR 1.2+X=9+.75X OR .25X=7.8 OR X=7.8/.25 OR X=31.2 GRAMS OF WATER MUST BE ADDED TO GET A SOLUTION OF 75% WATER & 25% IODINE.
PROOF 12*.1+31.2=.75(12+31.2) OR 1.2+31.2=.75(43.2) OR 32.4=32.4